The costs of doing business for a company can be found by adding fixed costs, such as rent, insurance, and wages, and variable costs, which are the costs to purchase the product you are selling. The portion of the company’s fixed costs allotted to this product is $300, and the supplier’s cost for a set of tile is $6 each. Let x represent the number of tile sets.
If b represents a fixed cost, what value would represent b?
Find the cost equation for the tile. Write your answer in the form C = mx + b.
The profit made from the sale of tiles is found by subtracting the costs from the revenue.
Find the Profit Equation by substituting your equations for R and C in the equation . Simplify the equation.
I have no idea where to begin. I REALLY need help. Please, I would appreciate any help that you could offer.

Answer by ankor@dixie-net.com(12706)

The costs of doing business for a company can be found by adding fixed costs, such as rent, insurance, and wages, and variable costs, which are the costs to purchase the product you are selling. The portion of the company’s fixed costs allotted to this product is $300, and the supplier’s cost for a set of tile is $6 each. Let x represent the number of tile sets.
:
If b represents a fixed cost, what value would represent b? "You can see from
the previous statement, that fixed cost is given as $300, therefore b = 300"
:
Find the cost equation for the tile. Write your answer in the form C = mx + b.
"The cost would be the cost for each tile, given as $6, times the number of
tiles which is x, plus the fixed cost, given as b," hence:
C = 6x + 30
:
The profit made from the sale of tiles is found by subtracting the costs from the revenue. " write this as Profit = R - C, substitute our cost equation for C
P = R - (6x + 30)
:
Find the Profit Equation by substituting your equations for R and C in the equation . Simplify the equation.
:
We don't know the revenue unless they tell us what the retail sale amt is for each tile.
For example if they sell each tile for $12. Revenue = 12x
then we can say: profit:
P = 12x - (6x + 300)
Simplify:
p = 12x - 6x - 300
p = 6x - 300
:
:
Did this help?